OpenStudy (anonymous):

Solve: 3(2x-1)+12=4x+1

OpenStudy (saifoo.khan):

x=-4

OpenStudy (anonymous):

6x-3+12=4x+1 .....2x=1+3-12 .....2x=-8 ....x=-4

OpenStudy (anonymous):

x=-4